Consider the initial value problem y – 2y – 8y = 0, y(0) = a, y(0) = 6 Find the value of a so that the solution to the in
Answer #1

gºs y-24 -8y=o y (o) = a. YO) = 6. the A.E of is given by, де - Duу йому. Чадам. 22-22 - 8 = 0. → 22-42428-8 = 0. -> r[ h-u)y (t) = (x-1) eat elett since t o y(t) 0. I ě 2 so as t s. - lin Y(t) = kila lent - & el has to be o. ef lim y(t) 20. =) gal=